Prove the solutions to this simple harmonic oscillation a=w^2x
1
Expert's answer
2020-03-05T09:36:34-0500

For this simple harmonic oscillation


"a=-\omega^2x"


"a=\frac{d^2x}{dt^2}"


"\frac{d^2x}{dt^2}=-\omega^2x"


Substitute "x(t)=e^{\lambda t}"


We have


"e^{\lambda t}(\omega^2+\lambda^2)=0"


"e^{\lambda t}\neq0"


"\lambda=i\omega" or "\lambda=-i\omega"


"x_1(t)=C_1e^{-i\omega t}" and "x_2(t)=C_2e^{i\omega t}"


The general solution is the sum of the above solutions


"x(t)=C_1e^{-i\omega t}+C_2e^{i\omega t}"


Using the Euler's identity we get


"x(t)=(C_1+C_2)\cos(\omega t)+i(-C_1+C_2)\sin(\omega t)"


In general


"x(t)=A_1\cos(\omega t)+A_2\sin(\omega t)"


or


"x(t)=A\cos(\omega t+\phi)"